OpenStudy (anonymous):

I really need help! MEDAL!!!! The speed of the current in a stream is 5 mi/hr. It takes a canoeist 300 minutes longer to paddle 5.5 miles upstream than to paddle the same distance downstream. What is the canoeist's rate in still water

OpenStudy (amistre64):

you know about vectors?

OpenStudy (amistre64):

canoe + stream is faster, but the same distance canoe - stream is slower, but the same distance since the distance remains the same, they have to equal each other out what is a good formula to find the distance covered if we know speed and time?

OpenStudy (anonymous):

No These are makeup assignments. So I'll need explanation

OpenStudy (anonymous):

And I honestly have no idea.

OpenStudy (amistre64):

this is a key part of the solution, youll need to find it before we move on. otherwise ill just feel like im doing all the work/thinking. we need to know how to find distance, using speed and time.

OpenStudy (amistre64):

if im running at 3 feet per sec, and i run for 5 sec ..... how far have i gone?

OpenStudy (anonymous):

distance=speed/time?

OpenStudy (amistre64):

soo close speed is measured in distance per time s = d/t, therefore d = st

OpenStudy (anonymous):

ohh, okay sorry

OpenStudy (amistre64):

3 feet/sec * 5sec = 15feet*sec/sec = 15feet

OpenStudy (amistre64):

anywhos we know the distance is the same sooo speed(up) * time(up) = speed(down) * time(down) this make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so it basically balances out

OpenStudy (amistre64):

yep up stream fights the current, so speed(up) = (canoe - river) down stream goes with the current, so speed(down) = (canoe + river) what do we know about the time frames?

OpenStudy (anonymous):

Well, the time frames are different? If that's a start

OpenStudy (anonymous):

should we subtract them?

OpenStudy (anonymous):

going 5.5 downstream takes 72 minutes going upstream its 372 minutes

OpenStudy (amistre64):

they are different :) takes 300 minutes longer to paddle upstream than downstream. time(up) = time(down) + 300

OpenStudy (amistre64):

soo plugging all this in: (canoe-river) (time(down)+300) = (canoe + river)(time(down)) river = 5 (c-5) (t+300) = (c+5)(t) any ideas from this?

OpenStudy (anonymous):

cancel the c-5 and the c+5?

OpenStudy (amistre64):

they dont cancel because they are not equal. and i just wrote up something that i couldnt verify so i deleted it. the distance traveled plays into this

OpenStudy (anonymous):

oh okay could you divide by t?

OpenStudy (anonymous):

and solve to get c?

OpenStudy (amistre64):

ugh, this isnt panning out for me, im missing something obvious :/ im thinking ... i need elimination for this? (c-5) (t+300) = d (c+5)(t) = d c(t+300)-5(t+300) = d ct +5t = d ct+300c-5t-1500 = d ct +5t = d ct+300c-5t = d+1500 ct +5t = d now we add ------------------------- ct+300c-5t = d+1500 ct +5t = d -------------------- 2ct+300c = 2d + 1500 not it

OpenStudy (amistre64):

im taking this to paper, typing is a bother

OpenStudy (anonymous):

I'm getting confused

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

i was seeing if my thought was getting somewhere, but its just running in circles :) let me organize it and see if we can approach this from the right angle is all

OpenStudy (anonymous):

okay, sounds great! thank you for helping by the way!

OpenStudy (amistre64):

speed = d/t c - r = d/(t+300) c + r = d/t elimination is the process where we remove one of the variables from the playing field; if we multiply the top equation by -1, and add the results we remove c. -c + r = -d/(t+300) c + r = d/t ----------------- 2r = d/t - d/(t+300) r and d are known so we can find t, right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

r = 5mi/60min d= 5.5 2(5/60) = 5.5(1/t - 1/(t+300)) can you tell me what t is?

OpenStudy (anonymous):

(5.5/t)-(5.5/t+300)????

OpenStudy (amistre64):

2(5/60) = 5.5(1/t - 1/(t+300)) 5/30 = 5.5( (t+300-t)/t(t+300) ) 1/6 = 5.5(300)/(t(t+300)) t(t+300) = 5.5(6)(300) t(t+300) = 5.5(6)(300) t^2 + 300t = 9900, which is a quadratic equation, not sure if youve covered these.

OpenStudy (anonymous):

We should go through it just for learnings sake

OpenStudy (amistre64):

maybe .... the shortcut is the quadratice formula given: ax^2 + bx + c = 0 x = [-b +- sqrt(b^2-4ac)]/(2a) in our case \[t=\frac{-300+\sqrt{300^2+4(9900)}}{2}\] or t=30

OpenStudy (anonymous):

so time is 30?

OpenStudy (amistre64):

time down stream is 30 minutes if ive kept my wits about me :)

OpenStudy (amistre64):

so, c+r = d/t, we know it all but c c = d/t - r c = 5.5/30 - 5/60

OpenStudy (anonymous):

c=6/60?

OpenStudy (amistre64):

now that sounds better, yes 6 miles per 60 min = 6 miles per hour

OpenStudy (anonymous):

so the answer is 6?

OpenStudy (amistre64):

with any luck, yes :)

OpenStudy (anonymous):

I'll check it really quick and tell you!

OpenStudy (amistre64):

and im sure ive taken a long way to get there, since the answer to this tends to be just a few short rows ...

OpenStudy (anonymous):

It's right!

OpenStudy (anonymous):

thank you for your help!

OpenStudy (amistre64):

youre welcome, if i can come up with the shorter solution process, ill post it lol

OpenStudy (anonymous):

haha okay, thanks again!

OpenStudy (amistre64):

using my first thought: \[(c-r)(t+300)=(c+r)(t)\] \[\frac{t+300}{t}=\frac{c+r}{c-r}\] \[1+\frac{300}{t}=\frac{c+r-r+r}{c-r}\] \[1+\frac{300}{t}=1+\frac{2r}{c-r}\] \[\frac{300}{t}=\frac{2r}{c-r}\] \[\frac{300}{t}=\frac{k(2r)}{k(c-r)}\] \[300=k\frac{2(5)}{60}=k/6\\ k=1800\\ t=1800(c-r) \] \[c+r=\frac{d}{1800(c-r)}\] \[c^2-r^2=\frac{d}{1800}\] \[c=\sqrt{r^2+\frac{d}{1800}}\] \[c=\sqrt{(5/60)^2+\frac{5.5}{1800}}=1/10\] 1 mile per 10 minutes, so 6 miles per 60 minutes 6 miles per hour

OpenStudy (amistre64):

it aint shorter, but at least we aint got to go all quadratic on it :) good luck